# A Practical Review of Mechanistic Interpretability for Transformer-Based Language Models

### v2

@@ 5.1.2 General Workflow for Open-ended Feature Study
The intermediate explanation generated in the Observe step is then interpreted by explainers to uncover features encoded in the representation. For instance, when a vocabulary projection method is used in the Observe step, a human explainer views the top tokens in the projected logit distribution and labels the feature implied by the tokens (e.g., if top tokens are all associated with arithmetic addition, then a human evaluator infers that the activation is encoding features relevant to additive operation in arithmetic [Rai & Yao 2024]). Similarly, for neuron and SAE activation visualization, a human explainer is instructed to examine the activation visualization for each neuron, and then indicate whether they have found a plausible theory to explain the activations. For instance, Elhage et al. (2022a) provided the following instructions to human explainer – “mark INTERPRETABLE if 80% or more of the strongest firings can be explained by a single rule or category (e.g. the word ‘apple’, or any phrase relating to music), and NOT INTERPRETABLE otherwise”. While the explainers are often humans, we can also automate this Explain step using machines, such as LLMs as explainers. This involves providing LLMs with intermediate explanations and prompting them to label the corresponding features, if applicable. For instance, Bills et al. (2023) provided intermediate explanations from both the vocabulary projection method and the neuron activation visualization to a GPT-4 model for feature labeling; the authors reported a strong correlation between human and GPT-4-generated explanations, supporting the feasibility of leveraging LLMs for feature explanation.
Figure 10: The interface prepared for the open-ended feature discovery using Neuronpedia Lin (2023). A human labeler or explainer uses this interface to label or explain the feature encoded by a neuron, based on the information presented within the interface.
@@ 5.2.1 General Workflow for Circuit Study
A circuit study typically involves the following steps: (1) Choose an LM behavior and a dataset: Select the LM behavior for which we are searching the circuit. (2) Define the computational graph: Describe the nodes and edges of the computational graph. (3) Localization: Identify all the important nodes and edges connecting them. (4) Interpretation: Explain the role of all nodes and edges in implementing the LM behavior. (5) Evaluation: Evaluate the faithfulness of the discovered circuit.
@@
In most existing literature Conmy et al. (2023); Syed et al. (2023); Hanna et al. (2024a; b), a circuit is defined as a directed acyclic graph (DAG) connecting individual computational components in the LM. For example, Hanna et al. (2024b) defined the GPT-2 model as DAG, where 144 attention heads (12 heads/layer $\times$ 12 layers) and the 12 FF sub-layers (1 per layer) in a GPT-2 model were considered nodes in its DAG and discovered circuits as sub-graphs from it. Following Elhage et al. (2021), many works also considered the connections between LM components in non-adjacent layers as valid edges in the LM’s DAG, due to the additive nature of the RS (Eq 1). That is, even if two components (e.g., the FF sub-layers in Layer 2 and Layer 5) are not connected through direct computations across adjacent layers, they are still connected in effect by writing to and reading from the RS, as shown in Figure 12. In practice, one can define a circuit at various levels of abstraction and granularity depending on the scope of interpretation. For example, instead of considering each LM component as a single node, one can define multiple nodes of the same LM component, one for each input position, when the position information may imply different roles of the same component Wang et al. (2022a); Nikankin et al. (2024), as shown in Figure 12.
Figure 12: (A) The IOI circuit, discovered by Conmy et al. (2023) as a reproduction of Wang et al. (2022a), where the computational graph of the model consists of nodes that are not directly connected across adjacent layers but still connected in effect by writing to and reading from the RS. (B) The arithmetic circuit, discovered by Nikankin et al. (2024), is represented as a computational graph where multiple nodes can be assigned to the same LM component, one per input position, when positional information suggests distinct functional roles for the same component.
@@
Once the computational graph for the LM is defined, the next step is to identify all the important nodes and edges responsible for implementing the behavior. This process is termed “localization”. Localization can be broken down into Localization of Nodes and Localization of Edges. (1) Localization of nodes: Intervention methods, such as zero ablation Olsson et al. (2022) and mean ablation Wang et al. (2022a), are commonly used to measure the importance of individual nodes in the model. These methods involve intervening in a node’s output and observing its impact on the model’s behavior, with important nodes being those whose intervention results in a deviation above a certain threshold in model performance. (2) Localization of edges: Path patching Wang et al. (2022a) is a commonly used intervention method to localize edges. Other intervention-based techniques, such as zero ablation, mean ablation, and resampling ablation, can also be applied for localization. However, these methods can be computationally-intensive as we need to measure the importance of each edge individually. To address this issue, automated localization techniques such as ACDC Conmy et al. (2023) and EAP Syed et al. (2023), have been proposed, as discussed in Section 4.2.3.
@@
Once all the important LM components and edges have been identified during the localization step, we interpret the functional role of each component. This interpretation step can be further divided into two sub-steps – (a) Generating a hypothesis: The first step involves generating a hypothesis about the function of each LM component. This process is often guided by analyzing the behavior of the component under relevant inputs using various techniques, such as visualization (e.g., attention visualization) and vocabulary projection methods. A human can then generate a plausible hypothesis based on the analysis (e.g., attention head H was observed to decrease the logit of the token it attends to and thus is hypothesized to function as a copy suppression head). (b) Validating the hypothesis: The hypothesis can be validated using various techniques, including both vocabulary projection and intervention-based methods. For instance, if the hypothesis suggests that attention head H is a copy suppression head, then knocking out (or noising) attention head H should remove the suppression effect. If the hypothesis is validated, the model component is considered interpreted; otherwise, further analysis and new hypotheses are needed.
@@ 5.3.1 General Workflow for Universality Study
The feature and circuit study can then be performed across LMs and/or tasks, as discussed in Sections 5.1.1 and 5.2.1.
Evaluation
Finally, we compare the similarities and differences of the discovered circuits and features across different LMs and/or tasks. Specifically, the features and circuits that are found across various LMs are considered to be universal, while those unique to specific LMs are considered model-specific. Likewise, one can draw conclusions about task-agnostic vs. task-specific circuits. It is important to note that the question of “do similar features and circuits exist across LMs and tasks” may not be binary; the experiments in practice instead reflect the degree to which universality holds. For instance, Gurnee et al. (2023) found that only 1-5% of the neurons were universal across five different GPT2-Small and GPT2-Medium models.
@@ 7.1 Feature Study
Earlier work studied neurons as a natural candidate for encoding features, where a high value at a particular position would indicate the presence of that feature. This led to the discovery of several interesting neurons such as sentiment neurons Tigges et al. (2023), Base64 neurons Elhage et al. (2022a), French neurons Gurnee et al. (2023), knowledge neurons Dai et al. (2022), skill neurons Wang et al. (2022b), positional neurons Voita et al. (2023), etc. The question of whether neurons encode specific features has been explored extensively, not only in the context of transformer-based LMs but also in earlier work on word embeddings Mikolov et al. (2013), recurrent neural networks (RNNs; Radford et al., 2017), and vision models Cammarata et al. (2020), predating the field of MI itself. However, many of these studies found that neurons within LMs are not monosemantic, i.e. they do not activate only for a single feature. In contrast, they are polysemantic, i.e. they activate in response to multiple unrelated features Elhage et al. (2022a); Gurnee et al. (2023); Elhage et al. (2022b). For instance, a single neuron could activate for both French texts and texts encoded in Base64. This polysemantic nature of neurons shows the initial hypothesis that each neuron encodes a single feature is not always true.
@@
The discovery of polysemantic neurons has led to the hypothesis of superposition, i.e., a model can represent a greater number of independent features than the number of available neurons Elhage et al. (2022b). In this framework, features are encoded as directions formed by a linear combination of multiple neurons rather than being tied to a single neuron. Although superposition enables the encoding of a larger number of features, it also introduces potential interference among overlapping representations, where activating one feature may unintentionally activate others that share the same representation space. Empirical findings suggest that the natural sparsity of feature activations helps mitigate this issue of interference Elhage et al. (2022b). Since only a small subset of features is typically active at any given time, the risk of unintended activations is reduced, allowing the neural model to maintain distinguishable representations despite the constraints imposed by superposition. Furthermore, models appear to prioritize encoding the most important features in more monosemantic fashion, while less critical features are either distributed across multiple neurons or omitted altogether, thereby minimizing performance degradation caused by interference Scherlis et al. (2022). Beyond its role in feature representation, superposition has also been observed to aid in data memorization for smaller datasets and in learning more generalized features for larger ones, with a notable transition between these regimes marked by the phenomenon of double descent Henighan et al. (2023).
@@
Recently, SAEs have gained popularity as a method for interpreting representations that encode features in superposition Bricken et al. (2023); Sharkey et al. (2022); Riggs (2023); Cunningham et al. (2023); Lieberum et al. (2024); Marks et al. (2024a); He et al. (2024). Various SAE variants with different architectures have been proposed, as discussed in Section 4.3. Early studies on SAEs have shown promising results, with the extracted features from SAEs appearing more interpretable than those from neurons, according to both human analysis and automatic explanation scores Bricken et al. (2023); Makelov (2024). Despite these promising findings, several challenges persist. For instance, SAEs do not perfectly reconstruct the activation which is often attributed to the trade-off between sparsity and reconstruction loss, raising concerns regarding the faithfulness of the extracted representations Gurnee (2024); Muhamed et al. (2024); Chanin et al. (2024); Anders & Bloom (2024). For instance, Templeton et al. (2024) speculated that SAEs may have only captured a fraction of features encoded in model activations, with many rare or highly specific concepts remaining undetected due to their infrequent activation. Similarly, recent studies Braun et al. (2024); Makelov et al. (2024) have also raised concerns regarding the functional usefulness of the extracted features, i.e., whether these features are actively used by the model for inference or merely reflect patterns present in the training data used to train the SAE. Although SAEs with improved architectures have been proposed to mitigate some of these limitations Muhamed et al. (2024); Braun et al. (2024), further advancements in SAEs are needed to ensure that the extracted features are both representative of the model’s underlying computation and practically useful for interpretability.

### v4

@@ 6.1.2 General Workflow for Open-ended Feature Study
The intermediate explanation generated in the Observe step is then interpreted by explainers to uncover features encoded in the representation. For instance, when a vocabulary projection method is used in the Observe step, a human explainer views the top tokens in the projected logit distribution and labels the feature implied by the tokens (e.g., if top tokens are all associated with arithmetic addition, then a human evaluator infers that the activation is encoding features relevant to additive operation in arithmetic [Rai & Yao 2024]). Similarly, for neuron and SAE activation visualization, a human explainer is instructed to examine the activation visualization for each neuron, and then indicate whether they have found a plausible theory to explain the activations. For instance, Elhage et al. (2022a) provided the following instructions to human explainer – “mark INTERPRETABLE if 80% or more of the strongest firings can be explained by a single rule or category (e.g. the word ‘apple’, or any phrase relating to music), and NOT INTERPRETABLE otherwise”. While the explainers are often humans, we can also automate this Explain step using machines, such as LLMs as explainers. This involves providing LLMs with intermediate explanations and prompting them to label the corresponding features, if applicable. For instance, Bills et al. (2023) provided intermediate explanations from both the vocabulary projection method and the neuron activation visualization to a GPT-4 model for feature labeling; the authors reported a strong correlation between human and GPT-4-generated explanations, supporting the feasibility of leveraging LLMs for feature explanation. In an interface like Neuronpedia Lin (2023), human evaluators and automatic LM explanations (i.e., the “auto-interp explanation” on the Neuronpedia interface in Figure 11) approach are combined to improve the efficiency of annotating feature explanations, as discussed in Section 6.1.3. Specifically, human evaluators are provided with LM-generated explanations as intermediate references alongside other intermediate explanations when formulating the final explanation. While this human–LM combination can accelerate the annotation process, it also risks biasing evaluators, who may become overly reliant on the LM’s output instead of producing independent and precise explanations.
6.1.3 Example Interfaces for Making Observations in Open-Ended Feature Study
We highlight two examples of interfaces from prior work that enable open-ended feature discovery: (1) Annotating neurons with features by Elhage et al. (2022a) (Figure 10), and (2) Feature interpretation using Neuronpedia Lin (2023) (Figure 11).
Figure 10: The neuron activation visualization developed by Elhage et al. (2022a), which labels neurons with features by analyzing the text on which they activate.
Annotating Neurons with Features
Elhage et al. (2022a) designed a simple interface to perform open-ended feature discovery, when they analyzed the FF neurons of a one-layer transformer model, as shown in Figure 10. The interface shows the intermediate explanations generated by neuron activation visualization (left) and vocabulary projection (right). Neuron activation visualization consists of highlighted text snippets that were sampled from a large corpus, focusing on paragraphs where the neuron exhibits the highest activation. We can see that the neuron seems to consistently activate on text encoded in base 64. Besides neuron activation visualization, the vocabulary projection method is applied to the same neuron to identify which output tokens it promotes or suppresses. The neuron seems to increase the logits of random mixed-case string tokens and suppresses common English words. This further supports the hypothesis that the neuron encodes base64-related features, as base64 strings often occur in sequence—making it sensible for the neuron to promote base64-like tokens as the next prediction. In summary, looking at the intermediate explanations from both the vocabulary projection method and neuron activation visualization, human evaluators can annotate the neuron to encode the base-64 feature.
Feature Interpretation using Neuronpedia
Figure 11: Example interface of Neuronpedia (Lin, 2023) for conducting open-ended feature studies. The example demonstrates a collection of intermediate explanations for interpreting an SAE feature (index 12,332) encoded in the residual stream activation at layer $20$ of the Gemma2-2b model. A human evaluator labels or explains the feature based on information present in this interface.
Neuronpedia (Lin, 2023) is a popular open-source platform for performing open-ended feature discovery. In particular, Neuronpedia serves as a general-purpose platform where researchers can upload collections of feature vectors discovered through different techniques (e.g., SAEs, custom probe vectors, etc.) for a given LM. The platform then generates an interactive dashboard for each feature vector by showcasing various intermediate explanations about the feature to support its open-ended feature study, facilitating both the Observe and the Explain steps. Figure 11 shows an example interface for conducting an open-ended feature study of residual stream activation at layer 2020 ($h^{20}_{i}$) of Gemma2-2b model Team et al. (2024). More specifically, the interface corresponds to a single SAE feature (index 12,33212{,}332) from the 16,00016{,}000 features obtained by projecting $h^{20}_{i}$ into a sparse representation ($20\text{-}\text{GEMMASCOPE}\text{-}\text{RES}\text{-}16k$) via an SAE. The interface consists of a collection of intermediate explanations, produced using various MI techniques, to guide human evaluators in the Explain step, as listed below:
  • Auto-Interp Explanation (upper-left panel): A concise natural language description (e.g., “French language text and dialogue” or “Program-related syntax and command structure”) generated by prompting LLMs of your choice (e.g, GPT-4 (Achiam et al., 2023)) with the top-activating text snippets for a given neuron or feature (Bills et al., 2023). Multiple, potentially unrelated, descriptions may be produced for the same SAE feature, as SAE features can sometimes be polysemantic. Accordingly, a quantitative measure (ranging from 0 to 1) that indicates how well the automatically generated explanations capture a feature’s behavior can be computed by employing another scoring LLM (“No Scores” currently shown in the demo interface). Following Bills et al. (2023), the scoring is performed by comparing the feature’s actual activation pattern with the activation predicted by the scoring LLM based on the auto-interp explanation. A higher score means the explanation consistently and accurately predicts the feature’s true activations, while a lower score suggests the explanation is incomplete or inaccurate.
  • Top Features by Cosine Similarity (middle-left panel): Features ranked based on their cosine similarity with the feature under investigation (i.e., feature of index 12,33212{,}332). The cosine similarity is calculated by comparing the target feature’s vector (i.e., the corresponding row vector of the SAE decoder, $W_{dec}[12332,:]\in\mathbb{R}^{d}$) with other features (i.e., $W_{dec}[j,:]\in\mathbb{R}^{d},j\neq 12332$).
  • Configuration (bottom-left panel): Configurations of the SAE architecture and the dashboard visualization, such as the dataset used to create the activation density plot and feature activation visualizations. The dataset can either be the original SAE training data or a different public dataset, such as pile-uncopyrighted (Gao et al., 2020) shown in the example interface. This flexibility is especially useful when the SAE training corpus is proprietary or confidential.
  • Statistical Information (upper-right panel): Upper-right panel consists of three intermediate explanations - vocabulary projection (left), activation density (top-right), and logit density histogram (bottom-right).
    • Vocabulary Projection (left): The feature vector $W_{dec}[12332,:]$ is projected to a vocabulary space or logit distribution by multiplying with $W_{U}$. The interface then displays the top-10 tokens with the highest and lowest logit scores to investigate the effect of the feature on the model output.
    • Activation Density (top-right): Histogram of randomly sampled non-zero activation values when provided with text input sampled from the dataset used to train SAEs or other public datasets (e.g., pile-uncopyrighted (Gao et al., 2020)).
    • Logit Density Histogram (bottom-right): Logit density histogram shows the logit distribution obtained from vocabulary projection of the feature vector. In the figure, we can observe that the feature vector assigns a negative logit score to most output tokens, which could indicate that their primary role could be the suppression of the set of output tokens.
  • Test activation with custom text (middle-right panel): An interactive form that allows practitioners to input custom text and observe whether the neuron activates. This helps validate the evaluator’s hypothesis. For example, if the feature is believed to detect French text, the evaluator can enter French and non-French texts to test whether the feature’s activation pattern aligns with the hypothesis.
  • Steer (middle-right panel): The interface includes a steering option that allows you to manually activate a feature with chosen activation strength and observe its influence on the model’s output. The Steer button is located next to the input form for testing activations with custom text, and it links to a separate interface designed for steering experiments.
  • Feature Activation Visualization (bottom-right panel): The visualization highlights tokens in input texts, randomly sampled from the dataset used to train SAE or other public datasets (e.g., pile-uncopyrighted (Gao et al., 2020)). that fall within specific activation intervals. To do this, activation values are divided into evenly spaced fractions of the feature’s maximum activation, and the corresponding texts that fall within each interval are extracted. This allows us to explore how the interpretability of a feature changes with varying activation strength.
Figure 12: Example interface designed by Bricken et al. (2023) for SAE feature interpretation, similar to the interface of Neuronpedia. However, the interface was designed mainly for demonstration purposes and is not a general-purpose SAE interpretation tool. It is not open-sourced either.
Now, based on these intermediate explanations, a human evaluator annotates the feature with a description in the Explain step. Neuronpedia already consists of completed Observe step for multiple open-weight models, including GPT-2 (Radford et al., 2019), Llama-2 (Touvron et al., 2023), and Gemma-2 (Lieberum et al., 2024), allowing users to begin directly from the Explain step. However, one can also upload new models along with trained SAE and other meta information (e.g., dataset for training SAE) to perform open-ended feature discovery on new models as well. Similar to Neuronpedia, Figure 12 is another example of an interface designed by Anthropic Bricken et al. (2023). It is created to perform an open-ended feature discovery on the SAE trained for the one-layer toy model analyzed in the paper. The interface presents overlapping intermediate explanations as Neuronpedia, though with fewer explanation types and slight differences in presentation. However, the interface is mainly designed for demonstration purposes. It is not open-sourced and is unavailable for use on new models by the broader research community.
@@ 6.2 Circuit Study
Techniques Advantages Limitations
Step 3: Localization
Zero, mean, and resampling ablation (Node/Edge localization) Faithful localization Computationally expensive and inefficient
Path patching (Edge localization) Faithful localization Computationally expensive and inefficient
ACDC (Auto Circuit DisCovery) Conmy et al. (2023) Faithful localization; Computationally efficient Computationally expensive
EAP Syed et al. (2023) Low computational requirements Less faithful than ACDC and activation patching
EAP-IG Hanna et al. (2024b) Faithful localization; Low computational requirements N/A
Step 4.1: Interpretation (Hypothesis Generation)
Visualization Intuitive; useful for rapid hypothesis generation about the function of the LM component Qualitative and easy to over-interpret; require follow-up causal tests
Vocabulary Projection Methods Provide important signals for what output tokens are promoted or suppressed by an LM component or activation Only decode features that can be represented in vocabulary space; early-layer projections can be unreliable without translators.
Step 4.2: Interpretation (Hypothesis Validation)
Vocabulary Projection Methods Quick verification on the effect of the LM component or activation on LM output Unreliable on early layers
Intervention-based Methods Causal evidence of necessity/sufficiency; quantify edge/node contributions. Out-of-distribution risks from corruption; compute-heavy at scale; require careful metric and corruption design
Probing Standard practice Require training data; correlational evidence only; detect feature presence but not usage
Table 8: Comparison of techniques for circuit study across Step 3 (Localization), Step 4.1 (Interpretation—Hypothesis Generation), and Step 4.2 (Interpretation—Hypothesis Validation).
@@ 6.2.1 General Workflow for Circuit Study
A circuit study typically involves the following steps: (1) Choose an LM behavior and a dataset: Select the LM behavior for which we are searching the circuit. (2) Define the computational graph: Describe the nodes and edges of the computational graph. (3) Localization: Identify all the important nodes and edges connecting them. (4) Interpretation: Explain the role of all nodes and edges in implementing the LM behavior. (5) Evaluation: Evaluate the faithfulness of the discovered circuit. Various MI techniques are used for conducting each step; we summarize their limitation and advantages in Table 8.
@@
(a) IOI circuit in GPT-2 Small discovered by Conmy et al. (2023).
(b) IOI circuit in GPT-2 Small discovered by Wang et al. (2022a).
Figure 14: Example illustrating computational graphs defined with different abstractions for the same IOI circuit. (a) The IOI circuit, discovered by Conmy et al. (2023), defines a computational graph to include LM components independent from input positions as nodes: mlml denotes the FF sublayer at layer $l$, $al.hj$ denotes the $j$-th attention head at layer $l$, and $final.inp$ refers to the logit score. One LM component (e.g., $a5.h5$) corresponds to only one node in the computational graph since they are position-independent. (b) The IOI circuit, discovered by Wang et al. (2022a), is represented as a computational graph where nodes corresponding to the same LM component at different input positions are considered distinct nodes, highlighting that one component may assume different functional roles depending on its positional context. Only attention heads are considered in this circuit, where $l.j$ denotes the $j$-th attention head at layer $l$.
@@
The second step of circuit discovery involves defining a computational graph of the model, which can be defined at different abstraction levels depending on the desired level of explanation detail for the model behavior. For example, Figure 14 shows two types of definition: Conmy et al. (2023) defined the computational graph such that a single component is represented by a single node across all token positions, whereas Wang et al. (2022a) treated the same component as different nodes, one for each decoding position. In addition, the definition of computational graphs can also vary in the granularity of their nodes. For example, Hanna et al. (2024b) defined the GPT-2 model as a computational graph, where 144 attention heads (12 heads/layer $\times$ 12 layers) and the 12 FF sub-layers (1 per layer) in a GPT-2 model were considered nodes in its computational graph and discovered circuits as sub-graphs from it. In contrast, Marks et al. (2024b) defined a computational graph of Pythia-70M Biderman et al. (2023) and Gemma-2-2B Team et al. (2024) model by considering the SAE features as nodes instead of attention heads, defining the graph at a more granular level. Beyond node definitions, an additional consideration in computational graphs lies in how edges are determined. Most existing literature Conmy et al. (2023); Syed et al. (2023); Hanna et al. (2024a; b) considers the connections between LM components in non-adjacent layers as valid edges in the LM’s computational graph, due to the additive nature of the RS (Eq 1). That is, even if two components (e.g., the FF sub-layers in Layer 2 and Layer 5) are not connected through direct computations across adjacent layers, they are still connected in effect by writing to and reading from the RS, as shown in Figure 14 (a).
@@
Once the computational graph for the LM is defined, the next step is to identify all the important nodes and edges responsible for implementing the behavior. This process is termed “localization”. Localization can be broken down into Localization of Nodes and Localization of Edges. (1) Localization of nodes: Intervention methods, such as zero ablation Olsson et al. (2022) and mean ablation Wang et al. (2022a), are commonly used to measure the importance of individual nodes in the model. These methods involve intervening in a node’s output and observing its impact on the model’s behavior, with important nodes being those whose intervention results in a deviation above a certain threshold in model performance. (2) Localization of edges: Path patching Wang et al. (2022a) is a commonly used intervention method to localize edges. Other intervention-based techniques, such as zero ablation, mean ablation, and resampling ablation, can also be applied for localization. However, these methods can be computationally-intensive as we need to measure the importance of each edge individually. To address this issue, automated localization techniques such as ACDC Conmy et al. (2023), EAP Syed et al. (2023), and EAP-IG Hanna et al. (2024b) have been proposed, as discussed in Section 5.2.3.
@@
Once all the important LM components and edges have been identified during the localization step, we interpret the functional role of each component. This interpretation step can be further divided into two sub-steps – (a) Generating a hypothesis: The first step involves generating a hypothesis about the function of each LM component. This process is often guided by analyzing the behavior of the component under relevant inputs using various techniques, such as visualization (e.g., attention visualization) and vocabulary projection methods. A human can then generate a plausible hypothesis based on the analysis (e.g., attention head $H$ was observed to decrease the logit of the token it attends to and thus is hypothesized to function as a copy suppression head). (b) Validating the hypothesis: The hypothesis can be validated using various techniques, including vocabulary projection, intervention-based methods, and probing. For instance, if the hypothesis suggests that attention head $H$ is a copy suppression head, then knocking out (or noising) attention head $H$ should remove the suppression effect. If the hypothesis is validated, the model component is considered interpreted; otherwise, further analysis and new hypotheses are needed.
@@ 6.3.1 General Workflow for Universality Study
The feature and circuit study can then be performed across LMs and/or tasks, as discussed in Sections 6.1.1 and 6.2.1.
Evaluation of Feature Universality
To measure similarity between features across different models, a common approach involves running each model on a common dataset (e.g., Pile test set (Gao et al., 2020)) and recording the activation patterns of individual features (i.e., neurons or SAE features) across all tokens in the dataset Gurnee et al. (2024). These activation patterns are treated as activation vectors representing each feature’s behavior. Similarity is then quantified by computing the Pearson correlation (e.g., Pairwise Pearson Correlation) between activation vectors from different models. For each feature in one model, the most similar counterpart in another model is identified by selecting the feature with the highest correlation. This process can be repeated across all features to construct a distribution of similarity scores. To account for sources of variation such as random initialization, architecture, or training noise, comparisons are often made against appropriate baselines or control settings. This method offers a principled way to identify and analyze features that are universal across models.
Evaluation of Circuit Universality
To determine whether a similar circuit is present across different models, we assess whether the models implement the same underlying algorithm for the behavior under study. Specifically, a circuit can be interpreted as an algorithm that is implemented by the components within the circuit. By interpreting the functional roles of these components, we can identify whether components with similar roles are present across models. Circuit universality is then evaluated by examining the overlap in functionally equivalent components across models under study Merullo et al. (2023).
It is important to note that the question of “do similar features and circuits exist across LMs and tasks” may not be binary; the experiments in practice instead reflect the degree to which universality holds. For instance, Gurnee et al. (2023) found that only 1-5% of the neurons were universal across five different GPT2-Small and GPT2-Medium models.
@@ 8.1 Feature Study
Neurons Description
Knowledge neurons Dai et al. (2022) FF neurons that store a relational fact (e.g., “The capital of Ireland is -> Dublin”); increasing or decreasing their activations directly affects whether the model recalls or suppresses the associated fact.
Skill neurons Wang et al. (2022b) A small set of FF neurons whose activation patterns are highly predictive of whether the model is performing a specific task (e.g., sentiment analysis); ablation of these neurons leads to a decrease in task performance.
Language-specific neurons Tang et al. (2024) Neurons that activate much more strongly for one or two languages compared to others; ablation of these neurons leads to a sharp drop in the model’s ability to understand and generate text in the language they encode.
Confidence regulation neurons Stolfo et al. (2024) Neurons in the final layer of LMs that adjust the model’s confidence (entropy of predictions) on what they are predicting, but don’t directly change which token is predicted.
Arithmetic neurons Rai & Yao (2024) Neurons that are crucial for arithmetic operations; ablating these neurons leads to a significant decrease in the performance of the arithmetic task.
Positional neurons Voita et al. (2024) Neurons that activate for specific token position ranges, i.e., they always activate for tokens within certain position ranges regardless of their meaning.
Dead neurons Voita et al. (2024) Neurons that never activate across a wide and diverse set of inputs, making them essentially unused by the model.
Safety-specific neurons Zhao et al. (2025) A small specialized set of neurons, found mainly in MHA layers, that are dedicated to recognizing and blocking harmful prompts (e.g., “how to build a bomb?”).
Table 9: A list of neurons identified in LMs, each associated with specific features or roles, along with their descriptions.
Earlier work studied neurons as a natural candidate for encoding features, where a high value at a particular position would indicate the presence of that feature. This led to the discovery of several interesting neurons such as sentiment neurons Tigges et al. (2023), knowledge neurons Dai et al. (2022), skill neurons Wang et al. (2022b), and positional neurons Voita et al. (2023), among others, as summarized in Table 9. The question of whether neurons encode specific features has been explored extensively, not only in the context of transformer-based LMs but also in earlier work on word embeddings Mikolov et al. (2013), recurrent neural networks (RNNs; Radford et al., 2017), and vision models Cammarata et al. (2020), predating the field of MI itself. However, many of these studies found that neurons within LMs are not monosemantic, i.e., they do not activate only for a single feature. In contrast, they are polysemantic, i.e., they activate in response to multiple unrelated features Elhage et al. (2022a); Gurnee et al. (2023); Elhage et al. (2022b). For instance, a single neuron could activate for both French texts and texts encoded in Base64. This polysemantic nature of neurons shows that the initial hypothesis that each neuron encodes a single feature is not always true.
@@
The discovery of polysemantic neurons has led to the hypothesis of superposition, i.e., a model can represent a greater number of independent features than the number of available neurons Elhage et al. (2022b). In this framework, features are encoded as directions formed by a linear combination of multiple neurons rather than being tied to a single neuron. Although superposition enables the encoding of a larger number of features, it also introduces potential interference among overlapping representations, where activating one feature may unintentionally activate others that share the same representation space. Empirical findings suggest that the natural sparsity of feature activations helps mitigate this issue of interference Elhage et al. (2022b). Since only a small subset of features is typically active at any given time, the risk of unintended activations is reduced, allowing the neural model to maintain distinguishable representations despite the constraints imposed by superposition. Furthermore, models appear to prioritize encoding the most important features in a more monosemantic fashion, while less critical features are either distributed across multiple neurons or omitted altogether, thereby minimizing performance degradation caused by interference Scherlis et al. (2022). Beyond its role in feature representation, superposition has also been observed to aid in data memorization for smaller datasets and in learning more generalized features for larger ones, with a notable transition between these regimes marked by the phenomenon of double descent Henighan et al. (2023).
@@
Recently, SAEs have gained popularity as a method for interpreting representations that encode features in superposition Bricken et al. (2023); Riggs (2023); Cunningham et al. (2023); Lieberum et al. (2024); Marks et al. (2024a); He et al. (2024). Various SAE variants with different architectures have been proposed, as discussed in Section 5.3. Early studies on SAEs have shown promising results, with the extracted features from SAEs appearing more interpretable than those from neurons, according to both human analysis and automatic explanation scores Bricken et al. (2023); Makelov (2024). Despite these promising findings, several challenges persist. For instance, SAEs do not perfectly reconstruct the activation which is often attributed to the trade-off between sparsity and reconstruction loss, raising concerns regarding the faithfulness of the extracted representations Gurnee (2024); Muhamed et al. (2024); Chanin et al. (2024); Anders & Bloom (2024). For instance, Templeton et al. (2024) speculated that SAEs may have only captured a fraction of features encoded in model activations, with many rare or highly specific concepts remaining undetected due to their infrequent activation. Similarly, recent studies Braun et al. (2024); Makelov et al. (2024) have also raised concerns regarding the functional usefulness of the extracted features, i.e., whether these features are actively used by the model for inference or merely reflect patterns present in the training data used to train the SAE. Although SAEs with improved architectures have been proposed to mitigate some of these limitations Muhamed et al. (2024); Braun et al. (2024), further advancements in SAEs are needed to ensure that the extracted features are both representative of the model’s underlying computation and practically useful for interpretability.